A circle whose centre lies on the line y=3x-1, touches both axes. Find the equation of the circle, which I found to be 4x^2+4y^2-4x-4y+1=0 and find the points where the circle intersects the line 2y+x=1.
Answer by ikleyn(52754)

.
A circle whose centre lies on the line y=3x-1, touches both axes. Find the equation of the circle,
which I found to be 4x^2+4y^2-4x-4y+1=0 and find the points where the circle intersects the line 2y+x=1.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~`

1. I confirm that you correctly determined the equation of the circle.


2. To find intersection points, express x = 1 - 2y from the equation of the straight line, and substitute it into the circle equation.

   You will get a single quadratic equation for x.

   Its solution will give you the x-coordinates of the intersection points.
   When you get x-coordinates, you can restore y-coordinates using the same equation 2y+x=1 of the straight line.